A man of mass 50kg ascends a flight of stairs 5m high in 5 seconds.If acceleration due to gravity is 10m/second-2.The power expended is???

1
Expert's answer
2022-03-24T11:01:27-0400

The power

"P=\frac{W}{t}=\frac{mgh}{t}"

"P=\frac{50*10*5}{5}=500\:\rm W"